OpenStudy (anonymous):

limit question. lim (x - > 5) 1/(x-5)^4

OpenStudy (anonymous):

\[\lim_{x \rightarrow 5} 1/(1+5)^4\]

OpenStudy (anonymous):

woops. the 1 is an x

OpenStudy (kirbykirby):

It would be infinity

OpenStudy (kirbykirby):

oh

OpenStudy (anonymous):

why so?

OpenStudy (kirbykirby):

Do you know L'Hopital's rule

OpenStudy (anonymous):

nope. is it the only way?

OpenStudy (anonymous):

the answer is infinity. just have no clue how to work it out.

OpenStudy (kirbykirby):

Ok what you can do is:

OpenStudy (kirbykirby):

wait i just noticed

OpenStudy (kirbykirby):

u mean it's x-5 in the denominator right? otherwise it wouldn't be infinity

OpenStudy (anonymous):

yes, x-5

OpenStudy (kirbykirby):

\[\lim_{x \rightarrow 5} \frac{x}{(x-5)^4}\]\[=\lim_{x \rightarrow 5} \frac{x}{(x-5)^4}\]\[=\lim_{x \rightarrow 5} \frac{x}{(x-5)^2(x-5)^2}\]\[=\lim_{x \rightarrow 5} \frac{x}{(x-5)(x-5)(x-5)(x-5)}\]\[=\lim_{x \rightarrow 5} \frac{x/x}{(x/x-5/x)(x/x-5/x)(x/x-5/x)(x/x-5/x)}\]\[=\lim_{x \rightarrow 5} \frac{1}{(1-5/x)(1-5/x)(1-5/x)(1-5)}\]

OpenStudy (kirbykirby):

oops the last factor should also be (1-5/x)

OpenStudy (anonymous):

awesome, thanks for your effort! just wondering line 6. why are you diving everything by x?

OpenStudy (kirbykirby):

I hope you see that the bottom factors are (1-1) when substituting x=5, so you get like "1/0" -> infinity

OpenStudy (kirbykirby):

hm it's true that was a useless step

OpenStudy (kirbykirby):

I was thinking that it was going to x->infinity at first, but i realized it was x->5 after lol

OpenStudy (anonymous):

ahhh, yes. i see that. thank you. is that a legal move though? because you are changing the function?

OpenStudy (anonymous):

for example, (x-5) doesn't equal (1-5/x)

OpenStudy (anonymous):

sorry, i guess it is. just would have never thought of that!

OpenStudy (kirbykirby):

Nope :) If you divide the top and bottom by x, it is perfectly valid (it's like you are dividing by "1") Example: \[\frac{2}{3} = \frac{\frac{2}{x}}{\frac{3}{x}}=\frac{2}{x}*\frac{x}{3}=\frac{2}{3}\]

OpenStudy (kirbykirby):

I divided the numerator by x as well: I did x/x = 1 in the numerator, and divided the whole denominator by x

OpenStudy (kirbykirby):

if you divide the numerator, you MUST also divide the denominator!

OpenStudy (anonymous):

awesome, this is just amazing. never would have figured out this.

OpenStudy (kirbykirby):

you will learn a neat theorem called "L'hopital's rule" later which will make the computation even easier :)

OpenStudy (anonymous):

k. thanks for your time. im now 'a fan' haha.

OpenStudy (kirbykirby):

Hehe awesome ;) Good luck with everything

OpenStudy (anonymous):

cheers